these problems go together.
DATA Mass of emply Beaker A Mass of Beaker A + grams of sodium carbonate Mass of empty Beaker Mass of Beaker B* grams of
Answer #1

2 Mass of sodium carbonate used = (69.0640 - 66.9790g. = 2.085g 3 Moles of NaCo₂ = 2.085g. 106 g/mol = 0.0197 mol, Mass of ca(4) Nazco + Cack Cacoz & R Nach, Mass of eacoz produced = {2.7600 - 1.005 g = 1.7559- Moles of eacon = 1:755g. looglimel - 8.